Grade 12Physical Chemistry

Electrolysis of a solution of HSO4- ions produces S2O82– . Assuming 75% current efficiency, what current should be employed to achieve a production rate of 1 mole of S2O8- - per hour ?

AKSHAY CHAUHAN
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
writting the reaction as:

2HSO4-...............+......>S2O82–+ 2H+.......+2e-
Let charge on elctron=q
from the equation we can see that 1 mole of S2O82- invoves two moles of electron.
So for 1 hour, the current should be = 2*NA*q/60*60 Ampere= I
but the efiiciency is 75% only, o to get this much actual amount the total current nedded=I’(let)
Hence 75% of I’=I
i.e., I’*75/100=I
Now please solve it.